Circular No. 8696

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2005nc AND GRB 050525A M. Della Valle, Istituto Nazionale di Astrofisica (INAF), Firenze; D.

Malesani, International School for Advanced Studies, Trieste; S. Benetti, INAF, Padova; G. Chincarini, University of Milano-Bicocca; L. Stella, INAF, Roma; and G. Tagliaferri, INAF, Brera, on behalf of a larger collaboration (Della Valle et al. 2006, Ap.J. Let., in press), report the discovery of an apparent supernova associated with the γ-ray burst GRB 050525A (Blustin et al. 2005, Ap.J. 637, 901). Observations obtained with the Telescopio Nazionale Galileo and with the Very Large Telescope (VLT) at the European Southern Observatory show that the early afterglow light curve is well described by a broken power law with a break 0.3 day after the γ-ray event. At ∼ 5 days after the burst, an apparent flattening is observed starting on May 30 (R magnitudes: May 30.28 UT, 23.96 ± 0.15; June 3.28, 24.03 ± 0.04; 7.24, 24.09 ± 0.07; 12.30, 24.01 ± 0.16), followed by a further dimming (June 30.15, R = 24.59 ± 0.09; July 6.19, 24.85 ± 0.08; Sept. 7.12, 25.04 ± 0.09; Oct. 1.00, > 25.1). A measurement of the position of SN 2005nc from the VLT frames yields α = 18^{h}32^{m}32^{s}.58, δ = +26^{o}20′22′′.6 (equinox 2000.0; uncertainty 0′′.25); Ryckoff et al. had given position end figures 32^{s}.6, 23′′.5 (uncertainty ∼ 1′′; cf. GCN 3468) for the afterglow. Inspection of a 2-hr integrated VLT spectrum (range 500–1000 nm, resolution 2 nm), obtained on 2005 June 28 (36 days after the GRB outburst, observer frame) confirms the redshift of the parent galaxy reported by Foley et al. (2005, GCN 3483) to be z = 0.606 and finds similarities with the early stages of the type-Ic broad-lined supernova 1998bw (Patat et al. 2001, Ap.J. 555, 900).

COMET P/2006 F1 (KOWALSKI) This comet (cf. IAUC 8690) has been found via additional astrometry,

including prediscovery images from Jan. 10, to be of short period; further astrometry and the following orbital elements appear on MPEC 2006-F49.

T = 2008 Feb. 6.3775 TT  ω = 184.^{o}7393
e = 0.122551  Ω = 125.0685 }2000.0
q = 4.134874 AU  i = 21.1773
a = 4.712383 AU n^{o} = 0.0963481 P = 10.23 years
2006 March 30  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2006 CBAT  Daniel W. E. Green